OpenStudy (anonymous):

Can someone please explain how i do this? 8b ---- A^ -5

OpenStudy (ranga):

8b / A^-5 = 8bA^5

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

How?

OpenStudy (ranga):

\[\Large \frac{ 1 }{ a^{-n} } = a^n\]

OpenStudy (ranga):

When you take the reciprocal the exponent will change sign.

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (ranga):

you are welcome.